How do I use substitution to get the answer x=0 and y=-4 for these problems: 10x+3y= -12
5x+4y= -16
I've tried quite a few times and keep getting fractions. Please help.

Answer by mananth(16949)

10 x + 3 y = -12 .............1
4 y = -16 -5 x ---------2
/ 4 =
y = -4 -1.25 x
Substitute y in (1)
10x+3(-4-1.25x )= -12
10x-12+ -3.75x = -12
10x-3.75x = -12 12
6.25 x = 0
/ 6.25
x= 0
Plug the value of x in (1)
10x+3= -12
10*0+3y = -12
0+3y=-12
3y=-12
/ 3
y= -4
m.ananth@hotmail.ca